OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

24 sec

OpenStudy (anonymous):

when h=0, it lands on the ground, so t(r-16t)=0 => t=0 & t=6 sec, which implies at time 0 & 6 the ball is on the ground, so ans is 6 sec

OpenStudy (anonymous):

so the answer is 24 seconds or 6 seconds?

OpenStudy (anonymous):

acceleration = -8 ,final velocity=0(at max.ht), initial velocity=96 => 0=96-8t =>t=12 , same time is taken for reaching the ground(neglecting air resistence) , so total time taken =24sec

OpenStudy (anonymous):

ans is 24 sure

OpenStudy (anonymous):

acceleration can't be -8 because for the half path the motion is against gravity & for the next half path the motion of the ball is towards earth so u can't use -8 for the whole path

OpenStudy (anonymous):

but time taken should be the same for both..isn't it

OpenStudy (anonymous):

in the equation u used before, v=u+at, final velocity will not be zero, when u are calculating the time for half the path, its max at highest pt. so u cant use this formula since u don't know the final velocity.

OpenStudy (anonymous):

final velocity must be zero at highest point.... review physics